In a multiple regression, the following sample regression equation is obtained: y-157 + 11.0x1 + 2.3x2 a. Predict y if x1 equals 17 and x2 equals 43. (Round your answer to 1 decimal place.) b. Interpret the slope coefficient of x1. As x1 increases by one unit, y is predicted to increase by 11.0 units, holding x2 constant. As x1 increases by one unit, y is predicted to increase by 2.3 units, holding x2 constant O As x1 increases by one unit, y is predicted to decrease by 2.3 units, holding x2 constant. As x1 increases by one unit, y is predicted to decrease by 11.0 units, holding x2 constant.
Answer #1

here,

y = 157 + 11x1 + 2.3x2

a) for x1 = 17, x2 = 43

y = 157 + 11*17 + 2.3*43 = 442.9

b) as x1 increases by one unit, y is predicted to increase by 11.0 units, holding x2 constant